OpenStudy (anonymous):

What equation represents the line that passes through the points (6, –3) and (–4, –9)?

OpenStudy (anonymous):

did u know formula for two point form if not then i will tell u?

OpenStudy (anonymous):

(x-x1)/(x-x2)=(y-y1)(y-y2) put the point and get equation

hero (hero):

Input each point into y = mx + b, creating two equations: -3 = 6m + b -9 = -4m + b Isolate b in each b = -3 - 6m b = 4m - 9 Set b = b, then isolate and solve for m: -3 - 6m = 4m - 9 9 - 3 = 6m + 4m 6 = 10m 6/10 = m 3/5 = m Substitute m into one of the formulas already isolated for b: b = 4m - 9 b = 4(3/5) - 9 b = 12/5 - 9 b = 12/5 - 45/5 b = (12 - 45)/9 b = -(45 - 12)/9 b = -32/9 Therefore: y = 3/5(m) - 32/9